OpenStudy (anonymous):

I don't get this at all!! Which describes the type of sequence in -99, -88, -66, -55....? a. both arithmetic and geometric b. geometric c. arithmetic d. neither arithmetic nor geometric Please explain as well! Thanks :)

jimthompson5910 (jim_thompson5910):

how do you test if a sequence is arithmetic?

OpenStudy (anonymous):

well i know that they don't have a constant pattern haha that's all i got :/

jimthompson5910 (jim_thompson5910):

You're right The more formal way to say this is that there isn't a common difference

jimthompson5910 (jim_thompson5910):

notice that the difference between -99 and -88 is 11, but the difference between -88 and -66 is 22

OpenStudy (anonymous):

so what's the answer then? I have nooo idea at all?? so it's not arithmetic... but is it geometric?? idk.. :O

jimthompson5910 (jim_thompson5910):

is there a common ratio?

OpenStudy (anonymous):

no i don't think so..

jimthompson5910 (jim_thompson5910):

we can test this by saying -55/(-66) = 5/6 but -66/(-88) = 3/4 So there isn't a common ratio

OpenStudy (anonymous):

sooo it's neither geometric OR arithmetic? I'm getting super confused,.. :(

jimthompson5910 (jim_thompson5910):

yes, it's "neither arithmetic nor geometric "

jimthompson5910 (jim_thompson5910):

so choice D

OpenStudy (anonymous):

yay thank you @jim_thompson5910 are you like really good at math or something??

jimthompson5910 (jim_thompson5910):

yes I am, you're very welcome

OpenStudy (anonymous):

Haha hi again @ParthKohli and thank you so much @jim_thompson5910 haha do you two know each other? lol and what's a harmonic sequence??

OpenStudy (anonymous):

I saw you were typing @ParthKohli but i didn't see anything pop up on my screen..did you reply?? or no? haha

jimthompson5910 (jim_thompson5910):

It appears s/he left, what's your question?

OpenStudy (anonymous):

oh haha i was wondering wat a harmonic sequence was but @ParthKohli said it wasn't a harmonic one hahhaa :)

jimthompson5910 (jim_thompson5910):

oh ok, yeah I was questioning that myself as well since the harmonic sequence involves fractions